OpenStudy (scott208):

If 2/3 p+ 6 =7/6 p, what is the value of p? Will fan and medal

OpenStudy (mjdennis):

Tag me when you come back online.

OpenStudy (scott208):

@mjdennis

Vocaloid (vocaloid):

hint: the common denominator of 2/3 and 7/6 is 6 so 2/3 p = 4/6 p

Vocaloid (vocaloid):

4/6 p + 6 = 7/6 p know where to go from here?

OpenStudy (scott208):

not really

OpenStudy (skullpatrol):